renalyn has a sheet of paper that is 20cm by 10 cm and makes a cylinder. What is the volume
Answer by mananth(16949)

Assume she rolls it on the 20 cm length
So 20 cm becomes thecicumference and 12 cm the height
circumference = 2pi r
2*pi*r=20
r= 20/2*pi
r=3.18cm
.
Volume of cylinder = pi*r^2*h
V= pi*3.18^2 *10
V= 317.7 cu.cm